OpenStudy (anonymous):

what times what equals 30 and adds up to -12

OpenStudy (anonymous):

1*30 2*15 3*10 5*6 Nothing.

OpenStudy (wolf1728):

A) x * y = 30 B) x + y = -12 x = 30/y 30/y + y = -12 30 + y² = -12y y² +12y +30 = 0 y = -3.5505 y = -8.4495 B) x -3.5505 = -12 x = -8.4495 x -8.4495 = -12 x = -3.5505 y = -3.5505 and x = -8.4495 -3.5505 -8.4495 = -12 -3.5505 * -8.4495 = 30